A pair of shoes and a sweater cost $98. If the sweater cost $16 more than the shoes, how much did the sweater cost?
Answer by JulietG(1812)

Z + s = 98
S = Z + 16
Therefore, Z + (Z+16) = 98
2Z + 16 = 98
2Z = 82
Z = 41
If the shoes cost 41, then the sweater costs 57
41 + 57 = 98